# A von Neumann algebraic approach to Quantum Theory on curved spacetime

### v3

@@ 7.2. Isotonic properties of $\{{\mathcal{M}}(\mathcal{O}):\mathcal{O}\in\mathcal{K}(M,g)\}$
We remind the reader that $\mathrm{ker}(\widetilde{G}_{k})=P_{\mathcal{O}_{k}}(\mathscr{D}(\mathcal{O}_{k},E))$. (See page 129 of [3].) So let $f=P_{\mathcal{O}_{2}}(g)$, where $g\in\mathscr{D}(\mathcal{O}_{2},E)$, be given. (Here $P_{\mathcal{O}}$ is the partial differential operator corresponding to the Klein-Gordon equation on $\mathcal{O}$.) The corollary will then follows from the preceding proposition once we notice that since it is the restriction of a function of compact support, the function $g|_{\mathcal{O}_{1}}\in C^{\infty}(\mathcal{O}_{1},E)$ must have bounded partial derivatives of all orders. ∎
@@
$U:f_{ext}+\overline{\mathrm{ker}(\widetilde{G}_{1})}\mapsto f+\overline{\mathrm{ker}(\widetilde{G}_{1})}\quad f\in\mathscr{D}(\mathcal{O}_{1},E)$
@@
$\int_{M}\langle(f-b_{f})_{ext},\upsilon\rangle,dV=\int_{M}\langle f-b_{f},\upsilon|_{\mathcal{O}_{1}}\rangle,dV=0.$
@@
By suitably rescaling the above unitary operator, we arrive at the following
@@
There is a unitary operator $v_{(2:1)}$ from $e_{(2:1)}H_{\mu_{2}}$ to $H_{\mu_{1}}$ which maps each $f_{ext}+\overline{\mathrm{ker}(\widetilde{G}_{2})}$, where $f\in\mathscr{D}(\mathcal{O}_{1},E)$, to $f+\overline{\mathrm{ker}(\widetilde{G}_{1})}$.
We now pass to comparing the algebras ${\mathcal{M}}(\mathcal{O}_{2},\mathcal{O}_{1})$ and ${\mathcal{M}}(\mathcal{O}_{1})$. By [30, Theorem 5.2.24] the GNS-spaces $H_{\omega_{k}}$ ($k=1,2$) are constructed from the action of the field operators on the Fock spaces $\mathcal{F}^{+}(H_{\widetilde{\mu}_{k}})$. We leave it as an exercise to verify that $e_{(2:1)}$ lifts to a projection $\Gamma(e_{(2:1)})$ from $\mathcal{F}^{+}(H_{\mu_{2}})$ onto $\mathcal{F}^{+}(e_{(2:1)}H_{\mu_{2}})$, and $v_{(2:1)}$ to a unitary operator $\Gamma(v_{(2:1)})$ from $\mathcal{F}^{+}(e_{(2:1)}H_{\mu_{2}})$ to $\mathcal{F}^{+}(H_{\mu_{1}})$, where $\Gamma(\cdot)$ stands for the second quantization map.
Let ${\mathcal{M}}(\mathcal{O}_{2}:\mathcal{O}_{1})$ denote the subalgebra of ${\mathcal{M}}(\mathcal{O}_{2})$ generated by the Weyl operators with symbols in $\mathrm{SOL}(\mathcal{O}_{2}:\mathcal{O}_{1})$ where $\mathrm{SOL}(\mathcal{O}_{2}:\mathcal{O}_{1})$ stands for the subset $\{\widetilde{G}_{2}(f_{ext}):f\in\mathscr{D}(\mathcal{O}_{1},E)\}$ of $\mathrm{SOL}(\mathcal{O}_{2})$. We claim that ${\mathcal{M}}(\mathcal{O}_{2}:\mathcal{O}_{1})$ commutes with $\Gamma(e_{(2:1)})$. The representation of $\mathrm{CCR}(\mathcal{O}_{2})$ on $\mathcal{F}^{+}(H_{\mu_{2}})$, therefore compresses to a representation of of $\mathrm{CCR}(\mathcal{O}_{2}:\mathcal{O}_{1})$ on $\mathcal{F}^{+}(e_{(2:1)}H_{\mu_{2}})$. To see this notice that for any $f\in\mathscr{D}(\mathcal{O}_{1},E)$ and $x_{1},\dots,x_{n}\in\mathrm{SOL}(\mathcal{O}_{2})$ we have that
$e_{(2:1)}^{n+1}(\widetilde{G}_{2}(f_{ext}))\otimes x_{1}\otimes\dots\otimes x_{n}$ $=$ $\widetilde{G}_{2}(f_{ext}))\otimes e_{(2:1)}(x_{1})\otimes\dots\otimes e_{(2:1)}(x_{n})$
$=$ $\widetilde{G}_{2}(f_{ext}))\otimes e_{(2:1)}^{n}(x_{1}\otimes\dots\otimes x_{n})$
and hence that $\Gamma(e_{(2:1)})a^{*}(\widetilde{G}_{2}(f_{ext}))=a^{*}(\widetilde{G}_{2}(f))\Gamma(e_{(2:1)})$. By duality we then similarly have that $\Gamma(e_{(2:1)})a(\widetilde{G}_{2}(f_{ext}))=a(\widetilde{G}_{2}(f_{ext}))\Gamma(e_{(2:1)})$, and hence that $\Gamma(e_{(2:1)})\phi(\widetilde{G}_{2}(f_{ext}))=\phi(\widetilde{G}_{2}(f_{ext}))\Gamma(e_{(2:1)})$. The Borel functional calculus now ensures that $\Gamma(e_{(2:1)})\pi_{\omega_{2}}(W(\widetilde{G}_{2}(f_{ext})))=\pi_{\omega_{2}}(W(\widetilde{G}_{2}(f_{ext})))\Gamma(e_{(2:1)}).$
Careful checking will similarly show that for any $f\in\mathscr{D}(\mathcal{O}_{1},E)$ we will in terms of $\Gamma(v_{(2:1)})$ have that $\Gamma(v_{(2:1)})a^{*}(\widetilde{G}_{2}(f_{ext}))=a^{*}(\widetilde{G}_{1}(f))\Gamma(v_{(2:1)})$ and hence that $\Gamma(v_{(2:1)})a^{*}(\widetilde{G}_{2}(f_{ext}))\Gamma(v_{(2:1)})^{*}=a^{*}(\widetilde{G}_{1}(f))$. Taking adjoints shows that we also have that $\Gamma(v_{(2:1)})a(\widetilde{G}_{2}(f_{ext}))\Gamma(v_{(2:1)})^{*}=a(\widetilde{G}_{1}(f))$, and so the same will be true of the field operators, namely that $\Gamma(v_{(2:1)})\phi(\widetilde{G}_{2}(f_{ext}))\Gamma(v_{(2:1)})^{*}=\phi(\widetilde{G}_{1}(f))$. So by the Borel functional calculus we must surely have that
$\Gamma(v_{(2:1)})\pi_{\omega_{2}}(W(\widetilde{G}_{2}(f_{ext})))\Gamma(v_{(2:1)})^{*}=\pi_{\omega_{1}}(W(\widetilde{G}_{1}(f))).$
The pair $({\mathcal{M}}(\mathcal{O}_{2},\mathcal{O}_{1}),\mathcal{F}^{+}(e_{(2:1)}H_{\mu_{2}}))$ is therefore unitarily isomorphic to $({\mathcal{M}}(\mathcal{O}_{1}),\mathcal{F}^{+}(H_{\mu_{1}}))$. That then yields the following observation:
@@
The pair $({\mathcal{M}}(\mathcal{O}_{2},\mathcal{O}_{1}),\mathcal{F}^{+}(e_{(2:1)}H_{\mu_{2}}))$ is unitarily isomorphic to the pair $({\mathcal{M}}(\mathcal{O}_{1}),\mathcal{F}^{+}(H_{\mu_{1}}))$.
7.3. Conclusion

### v4

@@ 7.3. Isotonic properties of $\{{\mathcal{M}}(\mathcal{O}):\mathcal{O}\in\mathcal{K}(M,g)\}$
We remind the reader that $\mathrm{ker}(\widetilde{G}_{k})=P_{\mathcal{O}_{k}}(\mathscr{D}(\mathcal{O}_{k},E))$. (See page 129 of [3].) So let $f=P_{\mathcal{O}_{2}}(g)$, where $g\in\mathscr{D}(\mathcal{O}_{2},E)$, be given. (Here $P_{\mathcal{O}}$ is the partial differential operator corresponding to the Klein-Gordon equation on $\mathcal{O}$.) The corollary will then follows from the preceding proposition once we notice that since it is the restriction of a function of compact support, the function $g{\upharpoonright}{\mathcal{O}_{1}}\in C^{\infty}(\mathcal{O}_{1},E)$ must have bounded partial derivatives of all orders. ∎
@@
$U:f_{ext}+\overline{\mathrm{ker}(\widetilde{G}_{2})}\mapsto f+\overline{\mathrm{ker}(\widetilde{G}_{1})}\quad f\in\mathscr{D}(\mathcal{O}_{1},E)$
@@
$\int_{M}\langle(f-b_{f})_{ext},\upsilon\rangle,dV=\int_{M}\langle f-b_{f},\upsilon{\upharpoonright}{\mathcal{O}_{1}}\rangle,dV=0.$
@@
In the following we shall write ${\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ for the subset $\{\widetilde{G}_{{\mathcal{O}}_{2}}(f_{ext}):f\in\mathscr{D}({\mathcal{O}}_{1},E)\}$ and similarly write ${\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ for the von Neumann subalgebra generated by the Weyl operators $\{\pi_{2}(W(x))\colon x\in{\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})\}$. By suitably rescaling the above unitary operator, we now arrive at the following
@@
There is a unitary operator $v_{(2:1)}$ from $e_{(2:1)}H_{\mu_{2}}$ to $H_{\mu_{1}}$ which maps each $\widetilde{G}_{{\mathcal{O}}_{2}}(f_{ext})$, where $f\in\mathscr{D}(\mathcal{O}_{1},E)$, to $\widetilde{G}_{{\mathcal{O}}_{1}}(f)$.
With the above technology now at our disposal we pass to an analysis of the von Neumann subalgebra of ${\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$. Given $f,g\in\mathscr{D}({\mathcal{O}}_{1},E)$, it follows fairly immediately that
$\tau_{2}(G_{2}(f_{ext}),G_{2}(g_{ext}))$ $=$ $\int_{{\mathcal{O}}_{2}}\langle f_{ext},G_{2}(g_{ext})\rangle\,dV$
$=$ $\int_{{\mathcal{O}}_{1}}\langle f,G_{1}(g)\rangle\,dV$
$=$ $\tau_{2}(G_{1}(f),G_{1}(g))$
The above is a fairly easy consequence of the definition of $G_{k}$ as described in [3, Proposition 3.5.1] and the fact that here $g$ is supported on $\mathcal{O}_{1}$. The map $T:{\mathrm{SOL}}({\mathcal{O}}_{1})\to{\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1}):G_{1}(f)\mapsto G_{2}(f_{ext})$ is therefore clearly a symplectic map. Writing ${\mathcal{A}}({\mathcal{O}}_{1})$, ${\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ for the corresponding Weyl algebras, there must by the Bogoliubov transform exist a *-isomorphism $\alpha_{(2:1)}$ from ${\mathcal{A}}({\mathcal{O}}_{1})$ onto ${\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ which maps Weyl operators of the form $W(G_{1}(f))$ ($f\in\mathscr{D}({\mathcal{O}}_{1},E)$) onto those of the form $W(G_{2}(f_{ext}))$. (Note that it clearly follows from [3, Corollary 4.2.11] that ${\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ may be regarded as a sub-algebra of ${\mathcal{A}}({\mathcal{O}}_{2})$.) Since by the preceding corollary we additionally have that
$\mu_{2}(G_{2}(f_{ext}),G_{2}(g_{ext}))=\mu_{1}(G_{1}(f),G_{1}(g))\mbox{ for all }f,g\in\mathscr{D}({\mathcal{O}}_{1},E),$
it is clear that the map $T$ also identifies the restriction of the two point function of $\omega_{{\mathcal{O}}_{2}}$ to ${\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$, namely
$(G_{2}(f_{ext}),G_{2}(g_{ext}))\to\mu_{2}(G_{2}(f_{ext}),G_{2}(g_{ext}))+\frac{1}{2}\tau_{2}(G_{2}(f_{ext}),G_{2}(g_{ext})),$
with the two point function for $\omega_{{\mathcal{O}}_{1}}$, namely
$(G_{1}(f),G_{1}(g))\to\mu_{1}(G_{1}(f),G_{1}(g))+\frac{1}{2}\tau_{1}(G_{1}(f),G_{1}(g).$
This then ensures that the state $\omega_{{\mathcal{O}}_{2}}\circ\alpha_{(2:1)}$ on ${\mathcal{A}}({\mathcal{O}}_{1})$ agrees $\omega_{{\mathcal{O}}_{1}}$. Here we made use of the fact that since $\alpha_{(2:1)}$ maps onto ${\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$, we are in the composition $\omega_{{\mathcal{O}}_{2}}\circ\alpha_{(2:1)}$ effectively dealing with the restriction of the state $\omega_{{\mathcal{O}}_{2}}$ to ${\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$. This leads to the following proposition:
@@
  1. (1)
    There exists a *-isomorphism $\alpha_{(2:1)}$ from ${\mathcal{A}}({\mathcal{O}}_{1})$ onto ${\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ which maps Weyl operators of the form $W(G_{1}(f))$ ($f\in\mathscr{D}({\mathcal{O}}_{1},E)$) onto those of the form $W(G_{2}(f_{ext}))$ and for which we additionally have that $(\omega_{2}{\upharpoonright}{{\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})})\circ\alpha_{(2:1)}=\omega_{{\mathcal{O}}_{1}}$.
  2. (2)
    Let $p_{(2:1)}$ be the projection from $H_{\omega_{2}}$ (the GNS Hilbert space of the pair $({\mathcal{M}}({\mathcal{O}}_{2}),\omega_{2})$ where ${\mathcal{M}}({\mathcal{O}}_{2})=\pi_{2}({\mathcal{A}}({\mathcal{O}}_{2}))^{\prime\prime}$) onto the Hilbert subspace generated by $\{\eta_{2}(W(x))\colon x\in{\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})\}$, and write ${\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ for the von Neumann subalgebra of ${\mathcal{M}}({\mathcal{O}}_{2})$ generated by the Weyl operators $\{\pi_{2}(W(x))\colon{\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})\}$. Then ${\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ commutes with $p_{(2:1)}$. In addition the GNS Hilbert space of the pair $({\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1}),\omega_{2}{\upharpoonright}{{\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})})$ may be identified with $p_{(2:1)}H_{\omega_{2}}$ and $\pi_{(2:1)}({\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1}))^{\prime\prime}$ with the restriction of ${\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ to $p_{(2:1)}H_{\omega_{2}}$ where in this instance $\pi_{(2:1)}$ denotes the *-isomorphism forming part of the GNS construction for the pair $({\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1}),\omega_{2}{\upharpoonright}{{\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})})$.
Proof.
The first claim is a direct consequence of the preceding discussion. The identification of $p_{(2:1)}H_{\omega_{2}}$ with the GNS Hilbert space of the pair $({\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1}),\omega_{2}{\upharpoonright}{{\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})})$ is a fairly straghtforward exercise. Once we show that ${\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ commutes with $p_{(2:1)}$, the final claim then follows by noting that strong convergence of nets in $\mathrm{span}\{\pi_{(2:1)}(W(x))\colon{\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})\}$ correspond to strong convergence of the matching nets in $\mathrm{span}\{\pi_{2}(W(x))\colon{\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})\}$. To see the claim regarding commutation with $p_{(2:1)}$ we firstly note that we will for any $x,y\in{\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ have that $\pi_{2}(W(x))\eta_{2}(W(y))=\eta_{2}(W(x)W(y))=e^{-i\tau_{2}(x,y)}\eta_{2}(W(x+y))$. Each $\pi_{2}(W(x))$ ($x\in{\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$) therefore maps $\overline{\mathrm{span}}\{\pi_{2}(W(x))\colon{\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})\}=p_{(2:1)}H_{\omega_{2}}$ back into itself. So for each $x\in{\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ and any $a,b\in H_{\omega_{2}}$ we have that
$\langle a,p_{(2:1)}\pi_{2}(W(x))p_{(2:1)}b\rangle=\langle a,\pi_{2}(W(x))p_{(2:1)}b\rangle$
and hence that $p_{(2:1)}\pi_{2}(W(x))p_{(2:1)}=\pi_{2}(W(x))p_{(2:1)}$. Since this is true for the generators of ${\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{2})$ it must be true for all elements of ${\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{2})$ with the claim regarding commutation then following by duality. ∎
It remains to show that the GNS standard form of the pair $({\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1}),\omega_{2}{\upharpoonright}{{\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})})$ as described above, embeds into the GNS standard form of the pair $({\mathcal{M}}({\mathcal{O}}_{2}),\omega_{2})$. For this task we fisrtly need the following specialization of [32, Theorem 5.2.24(b)(i)-(iii)].
Lemma 7.9.
Let $p_{(2:1)}$ and $H_{\omega_{2}}$ be as before. Then the following holds true:
  1. (1)
    With $\widetilde{e}_{(2:1)}$ denoting the projection of $\mathcal{H}_{\mu_{2}}$ onto the Hilbert subspace of $\mathcal{H}_{\mu_{2}}$ generated by ${\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$, the space $p_{(2:1)}H_{\omega_{2}}$ is the bosonic (symmetrized) Fock space $\mathcal{F}^{+}(\widetilde{e}_{(2:1)}\mathcal{H}_{\mu_{2}})=\Gamma(\widetilde{e}_{(2:1)})\mathcal{F}_{+}(\mathcal{H}_{\mu_{2}})$. In particular $p_{(2:1)}=\Gamma(\widetilde{e}_{(2:1)})$ and $H_{\omega_{2}}=\mathcal{F}_{+}(\mathcal{H}_{\mu_{2}})$.
  2. (2)
    With $\Psi_{2}$ denoting the vacuum vector of the Fock space $\mathcal{F}_{+}(\mathcal{H}_{\mu_{2}})$, the subspace of the finite complex linear combinations of $\Psi_{2}$ and all vectors of the form
    $a^{*}(x_{1})\dots a^{*}(x_{n})\Psi_{2}\mbox{ for }n=1,2,\dots$
    where $x_{k}\in{\mathrm{SOL}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ for all $1\leq k\leq n$, is a dense subspace of $p_{(2:1)}H_{\omega_{2}}$.
Proof.
The proof follows from a straightforward modification of the first paragraph of the proof of [33, Lemma A.2]. ∎
The above lemma now enables us to note the following:
Lemma 7.10.
Let $p_{(2:1)}$ and $H_{\omega_{2}}$ be as before and consider the GNS standard form $({\mathcal{M}}({\mathcal{O}}_{2}),H_{\omega_{2}},J_{{\mathcal{O}}_{2}},\mathscr{P}_{{\mathcal{O}}_{2}})$. Then $p_{(2:1)}H_{\omega_{2}}$ is an invariant subspace of the modular conjugation operator $J_{{\mathcal{O}}_{2}}$.
Proof.
Taking into account the unitary equivalence of the Araki-Woods and Kay-Wald approached noted in subsection 7.1, the claim follows from the preceding lemma considered alongside the final two equalities in [14, Theorem 40(3)]. See also the fragment of the proof of [14, Theorem 40(6)] showing that $J_{s}$ preserves the space $\mathcal{H}$ defined there. ∎
We finally arrive at the main result for this subsection.
Proposition 7.11.
There is a standard form homomorphism from the standard form of $({\mathcal{M}}({\mathcal{O}}_{1}),\omega_{{\mathcal{O}}_{1}})$ into the standard form of $({\mathcal{M}}({\mathcal{O}}_{2}),\omega({\mathcal{O}}_{2}))$.
Proof.
Since by Proposition 7.8 the pairs $({\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1}),\omega_{2}{\upharpoonright}{{\mathcal{A}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})}$ and $({\mathcal{A}}({\mathcal{O}}_{1}),\omega_{{\mathcal{O}}_{1}})$ are *-isomorphically equivalent, it is clear that they will generate ismorphic standard forms for ${\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})$ and ${\mathcal{M}}({\mathcal{O}}_{1})$ respectively. It therefore remains to show that the standard form of the pair $({\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1}),\omega_{2}{\upharpoonright}{{\mathcal{M}}({\mathcal{O}}_{2}:{\mathcal{O}}_{1})}$ embeds into the standard form of the pair $({\mathcal{M}}({\mathcal{O}}_{2}),\omega_{2})$. This in turn follows from a combination of part (2) of Proposition 7.8, and an application of Lemma 3.6 to what we noted in Lemma 7.10. ∎
7.4. Emergent framework